OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 4 points) Which one of the following does not find the value of f ′(3) for f(x) = 3x^2 – 4x + 1? A) f'(3)=lim as h approaches 0 of 3(3+h)^2-4(3+h)+1-(3(3)^2-4(3)+1)/h B)f'(3)=lim as h approaches 0 of 3(3+h)^2-4(3+h)+1-3(3)^2+4(3)-1)/h C)f'(3)= lim as x approaches 3 of 3x^2-4x+1-16/x-3 D)All of these find the value for f'(3)

OpenStudy (anonymous):

@ganeshie8 :) Please helpp!! will give medal to best answer!!!!!!

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@jim_thompson5910

ganeshie8 (ganeshie8):

you need to know below two limit definitions of derivative :\[f'(a) = \lim\limits_{h\to 0}~\dfrac{f(a+h)-f(a)}{h} = \lim\limits_{x\to a}~\dfrac{f(x) - f(a)}{x-a}\]